Circular No. 6907

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cc IN NGC 5172 W. D. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University

of California at Berkeley, report their discovery of an apparent supernova (mag ∼ 18.1) during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The object was found and confirmed on unfiltered CCD images obtained on May 15.3 and 16.3 UT, respectively. SN 1998cc is located at α = 13^{h}29^{m}19^{s}.31, δ = +17^{o}02′42′′.4 (equinox 2000.0), which is ∼ 24′′.5 due south of the nucleus of NGC 5172. No stars at the position of the supernova are visible in the Digital Sky Survey image of the field, while KAIT images obtained on Apr. 22 and 28 show a very faint source of roughly constant magnitude. Whether this faint source is an underlying H II region or the very early phase of the supernova remains to be determined by spectroscopic observations.

S. Jha, P. Garnavich, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram (range 350–760 nm) was taken of SN 1998cc by P. Berlind at the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on May 16.3 UT. The spectrum displays He I and strong Fe II absorption, indicating that the supernova is of type-Ib before maximum light. The host galaxy recession velocity, determined from narrow Hα emission, is 4080 km/s.

SUPERNOVA 1998bu IN NGC 3368 Y. D. Mayya and I. Puerari, Instituto Nacional de Astrofísica Optica

y Electrónica, Puebla; and O. Kuhn, Instituto de Astronomía, UNAM, Ensenada, report the following photometric magnitudes (± 0.05 mag) obtained with the Observatorio Astronómico Nacional 2.1-m telescope (+ CAMILA/NICMOS 3 camera) at San Pedro Martir, Mexico: May 11.30– 11.31 UT, J = 11.77, H = 11.73, K = 11.81; 12.15–12.19, 11.67, 11.66, 11.67; 14.15–14.21, 11.49, 11.58, 11.42.

COMET C/1998 J1 (SOHO) F. Farrell, Christies Beach, S. Australia, reports that on May 16.36

UT, 7×50 binoculars revealed a tail 3^{o} long in p.a. 100^{o} (m_{1} = 2.4, coma diameter ∼ 5′), and that the comet was faintly visible to the naked eye (though the tail was not).

1998 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green